# Attention Projection Mixing and Exogenous Anchors

### v1

@@ Abstract
[…]
@@ 1 Introduction
A primary technique for stabilizing deep networks is the use of residual connections. In Transformers, a layer’s input is added to its output, preserving a pathway for gradients and early-layer information. Despite this, studies show that token-level information can become diluted in deeper layers, a phenomenon linked to over-smoothing (Shi et al., 2022; Zhou et al., 2021). This has spurred interest in more direct mechanisms for preserving early representations.
@@
(a) For NuResFormer, the anchor components are sourced from the first layer. For ExoFormer, they are provided by dedicated exogenous projections. All residualized sources are RMSNormed before being mixed with the current layer’s projections, and the scaled dot product attention (SDPA) output is gated.
(b) Training curves showing validation loss across steps. All models have similar parameter counts (453M–454M) except ExoFormer, which uses a dedicated exogenous anchor and has 457M parameters. ’Naive Combination’ refers to the unmodified addition of Gated Attention and ResFormer.
Figure 1: Overview of the proposed architectures and their training performance.
We first develop a unified framework for cross-layer mixing across all attention pathways: queries, keys, values, and gate logits. A key insight is that applying RMSNorm to the residual sources before mixing resolves distributional mismatch, enabling stable and beneficial reuse. The internal-anchor version, which we term NuResFormer (Normalized unified), serves as a comparative foundation.
@@
Our primary contribution is ExoFormer, which resolves this tension by learning a set of dedicated exogenous anchor projections outside the sequential layer stack. This decoupling proves consistently beneficial: every ExoFormer variant outperforms its corresponding NuResFormer counterpart in perplexity while remaining competitive in downstream accuracy despite no increase in width or depth, with the dynamic variant achieving the best results. Furthermore, ExoFormer exhibits a paradoxical combination of high performance and signs of representation collapse. We explain this via an Offloading Hypothesis: the external anchor assumes the role of preserving token identity, allowing the sequential layers to specialize almost exclusively in refinement.
@@ 2 Related Work
Most directly relevant are ResFormer and Gated Attention. Zhou et al. (2025) introduced Value Residual Learning (ResFormer), which adds a residual connection from the first layer’s value vectors ($V_{1}$) to the value projections of all subsequent layers. This simple yet effective method was shown to greatly improve model performance and data efficiency, highlighting the benefit of explicitly reusing early content representations. However, preliminary attempts to residualize queries and keys were found to be unstable. Concurrently, gated attention mechanisms have been explored to introduce dynamic, input-dependent modulation to the attention output, improving expressiveness and training stability (Qiu et al., 2025). Our work generalizes residual learning beyond values and integrates it with gating via normalized mixing.
@@
Recent work has sought to improve cross-layer information flow in Transformers beyond simple residual connections. Zhu et al. (2025) proposed Hyper-Connections that expand residual stream width. More recently, Xie et al. (2026) introduced Manifold-Constrained Hyper-Connections (mHC), restoring an identity-like signal-preservation property to hyper-connected architecture.
@@
While both MUDDFormer and our work focus on enhancing the attention pathways, they operate at different levels and can coexist. MUDDFormer primarily addresses the input to the QKV projections. In contrast, our unified mixing framework focuses on mixing the projected QKV tensors themselves. Our approach can be seen as a form of residual blending after projection, whereas MUDDFormer enriches the input before projection.
@@ 3.1 Preliminaries and Notation
Granularity of learned coefficients.
We consider three coefficient granularities throughout for $\lambda$:
  • Scalar (S): a single coefficient shared across all channels.
  • Headwise (H): one coefficient per head, broadcast across the head dimension $d_{k}$.
  • Elementwise (E): one coefficient per channel.
@@ 3.2 Multi-Head Attention
$A_{n}=\mathrm{softmax}\left(\frac{Q_{n}K_{n}^{\top}}{\sqrt{d_{k}}}\right)\in\mathbb{R}^{T\times T},\qquad U_{n}=A_{n}V_{n}\in\mathbb{R}^{T\times d_{\text{model}}},$ (2)
@@
In NuResFormer/ExoFormer, the above stages use mixed tensors $\widehat{Q}_{n},\widehat{K}_{n},\widehat{V}_{n}$ from Eq. (10). Note that while heads are not shown explicitly in this formulation, the practical implementation follows the standard multi-head mechanism.
@@ 3.3 Gated Attention
In our attention block, we apply gating to the concatenated multi-head output (i.e., after Stage 4 and before Stage 5). Let $G_{n}\in\mathbb{R}^{T\times(hd_{k})}$ denote the gate logits, computed from $H_{n-1}$:
@@ 3.4 Unified Mixing Formulation
We now present a general framework for mixing. The core idea is to enrich each layer’s attention pathways with a set of persistent anchor projections $\{Q_{\text{anc}},K_{\text{anc}},V_{\text{anc}},G_{\text{anc}}\}$ that are reused, via learnable mixing, across all layers.
@@
The anchor projections are a fixed set of tensors defined once for the entire model. In our work, we explore two instantiations:
@@
    $Q_{\text{anc}}=H_{0}W^{Q}_{\text{anc}},\;K_{\text{anc}}=H_{0}W^{K}_{\text{anc}},\;V_{\text{anc}}=H_{0}W^{V}_{\text{anc}},\;G_{\text{anc}}=H_{0}W^{G}_{\text{anc}},$ (9)
    with independent weight matrices $W^{Q}_{\text{anc}},W^{K}_{\text{anc}},W^{V}_{\text{anc}},W^{G}_{\text{anc}}$.
@@
$\widehat{S}_{n}=\lambda^{S}_{n,1}\odot\mathrm{RMSNorm}(S_{\text{anc}})+\lambda^{S}_{n,2}\odot S_{n},\qquad\forall S\in\{Q,K,V,G\},$ (10)
@@ 3.5 First Layer Tension and ExoFormer
While these roles appear misaligned—universal anchors favor invariance, while progressive computation necessitates change—they can theoretically coexist because it is an optional pathway modulated by learned mixing coefficients ($\lambda_{n,1},\lambda_{n,2}$). However, this forced coexistence compels the first layer to make compromises, inherently limiting its effectiveness in both roles.
We explored strategies to decouple the “anchor” role from the first attention layer. Our most successful approach, termed ExoFormer (where “Exo” signifies an external, outside-the-system anchor), instantiates the general framework of Section 3.4 with the dedicated exogenous projections of Eq. (9).
This decoupling proves beneficial: in our experiments, every ExoFormer variant outperformed its corresponding NuResFormer counterpart in perplexity, despite adding only four additional projection matrices with no change to depth or width. The result confirms that separating the reusable-anchor task from the progressive-computation task allows each component to specialize, leading to better downstream performance.
@@ 3.6 Dynamic Mixing (DM) Module
Building upon the unified formulation, we take inspiration from MUDDFormer’s Depth-wise Aggregate (DA) module (Xiao et al., 2025) and introduce a dynamic variant where the learnable parameters are modulated by context-dependent scaling factors computed from the layer input $H_{n-1}$ using a small MLP. This allows the model to adapt its mixing strategy based on the specific context. It is possible to combine both modules (DA and DM) for increased efficiency since they share the same structure.
@@
For each component $S\in\{Q,K,V,G\}$:
$\widehat{S}_{n}=(\lambda^{S}_{n,1}\odot\gamma^{S}_{n,1})\odot\text{RMSNorm}(S_{\text{anc}})+(\lambda^{S}_{n,2}\odot\gamma^{S}_{n,2})\odot S_{n}$ (12)
@@ 4.1 Experimental Setups
All models use a pre-normalized modern Transformer architecture with SwiGLU activations (Shazeer, 2020), QKNorm (Henry et al., 2020), and rotary position embeddings (Su et al., 2023). We follow prior work by initializing projection and classification layers to zero (Yang et al., 2022), removing bias terms (except for the dynamic mixing module) (Chowdhery et al., 2022), applying z-loss regularization (debrébisson2016zlossshiftscaleinvariant), and disabling dropout. Training is performed on a 10B-token subset of the FineWeb-Edu dataset (Penedo et al., 2024), which aligns with compute-optimal training as informed by Chinchilla scaling laws (Hoffmann et al., 2022).
Model ARC-c ARC-e HellaSwag OBQA PIQA Wino Avg. Acc PPL Params (M)
Base Transformer 30.97 63.38 42.02 33.60 67.30 51.54 48.14 14.79 454
Gated Attention 30.89 64.69 42.73 34.00 67.14 53.35 48.80 14.64 453
ResFormer (Value Residual) 33.62 64.86 43.49 34.40 68.88 52.64 49.65 14.32 454
Naïve Combination (ResFormer+Gate) 32.94 64.52 43.47 32.60 68.34 52.64 49.09 14.25 453
NuResFormer (Internal Anchor) 33.70 65.07 44.23 33.60 68.34 53.12 49.68 14.15 453
ExoFormer (no normalization on anchor) 32.08 63.76 43.49 34.40 68.77 55.64 49.69 14.30 457
Dynamic ExoFormer 33.36 65.87 44.54 34.40 68.28 55.17 50.27 14.09 457
ExoFormer 34.73 64.65 44.28 36.40 67.74 53.59 49.85 14.13 457
Table 1: Performance comparison of model variants on 6 multiple-choice downstream tasks. Reported metrics include accuracy (%), validation loss, perplexity (PPL), and parameter counts. Dynamic ExoFormer achieves the highest overall accuracy and the lowest perplexity, demonstrating the advantages of decoupling the anchor from the first layer and using context-aware mixing. All models use full normalization on anchors unless noted otherwise, and NuResFormer and ExoFormer employ elementwise mixing coefficients. A detailed analysis of mixing coefficient granularity is provided in Appendix B.
We optimize 2D parameters with Muon (specifically the Polar Express variant; Amsel et al. (2025)), applying a cautious weight decay of 0.1 (Chen et al., 2025), while 1D parameters are trained using AdamW without weight decay. Muon was utilized because it represents a state-of-the-art optimizer for large-scale LLM training (Liu et al., 2025). Gradient norms are clipped to 1.0, and all models follow the same optimization setup, to ensure fair comparison across architectures. Training uses a global batch size of 262,144 tokens, a sequence length of 2,048, and a total budget of 10 billion tokens.
@@ 4.2 The Role of Anchor Normalization
As a mild isotropization operator, RMSNorm projects representations onto the unit sphere while preserving directional information and removing scale differences. The unit vector retains the relative alignment between dimensions, which encodes semantic and syntactic features. Second, RMSNorm as a scale-dependent rational function introduces a non-linear transformation into the otherwise linear residual pathway.
This is empirically supported by an analysis of the learned mixing coefficients in unnormalized models. For instance, in the ExoFormer variant without residual normalization, the proportion of near-zero coefficients (below 0.001) for $\lambda_{1}$ (the strength of anchor signal) is approximately triple that of the normalized variant. This suppression of the anchor pathway suggests the model is actively compensating for distributional mismatch.
We hypothesize that the naïve combination underperforms for the above reason. When applied to such unbalanced mixtures, $\sigma(G)$ must compensate for distributional mismatch rather than execute content-aware filtering, resulting in unstable oscillations during training (see Figure 1(b)) and worse downstream accuracy than ResFormer in isolation.
@@ 4.3 Extending Mixing to Q, K, and G Pathways
We systematically evaluate the contribution of adding residual pathways for queries ($Q$), keys ($K$), and gating logits ($G$), in addition to the established value ($V$) residual. Table 3 in the Appendix reports the performance of models that residualize different subsets of $\{Q,K,V,G\}$, with and without applying RMSNorm to the residual sources.
@@
Consistent with prior observations (Zhou et al., 2025), we find a clear hierarchy of stability. Models attempting to use unnormalized Q/K residuals without QKNorm proved impossible to optimize, exhibiting divergent loss in preliminary runs. Critically, adding QKNorm alone stabilizes training and recovers baseline performance. We hypothesize this is because QKNorm compresses the scale of $Q$ and $K$, mitigating the distributional mismatch that arises when injecting early-layer routing signals into deeper layers. Furthermore, adding explicit RMSNorm to the Q/K residual sources on top of QKNorm yields the best results, enabling positive reuse.

### v2

# Attention Projection Mixing with Exogenous Anchors

@@ Abstract
Cross-layer reuse of early attention projections can improve optimization and data efficiency, but it creates a structural conflict: the first layer must simultaneously act as a stable, reusable anchor for all deeper layers and as an effective computational block. We show this ”first-layer tension” is a hidden limiter of internal-anchor designs. We propose ExoFormer, which resolves the conflict by learning exogenous anchor projections outside the sequential layer stack, decoupling the anchor role from computational refinement. We introduce a unified normalized mixing framework that mixes queries, keys, values, and gate logits using learnable coefficients (exploring coefficient granularities: elementwise/headwise/scalar), and we show that normalizing anchor sources is key to stable reuse. ExoFormer variants consistently outperform their internal-anchor counterparts, and the dynamic variant yields $\sim\!1.5$ downstream accuracy points while matching validation loss using $\sim\!1.5\times$ fewer tokens than Gated Attention. We explain this efficacy via an Offloading Hypothesis: external anchors preserve essential token identity, allowing layers to specialize exclusively in refinement. We release code and models to facilitate future research.
Machine Learning, ICML, Transformers, NLP
@@ 1 Introduction
Studies show that token-level information can become diluted in deeper layers, a phenomenon linked to over-smoothing (Shi et al., 2022; Zhou et al., 2021). This has spurred interest in more direct mechanisms for preserving early representations.
@@
Figure 1: For NuResFormer, the anchor components are sourced from the first layer. For ExoFormer, they are provided by dedicated exogenous projections. All residualized sources are normalized before being mixed with the current layer’s projections, and the scaled dot product attention (SDPA) output is gated.
We introduce a unified framework for cross-layer mixing across all attention pathways through systematic evaluation of the contribution of mixing queries ($Q$), keys ($K$), and gating logits ($G$), in addition to the established value ($V$) residual using different coefficient granularities (elementwise, headwise, scalar). A key insight is that applying RMSNorm to the residual sources before mixing resolves distributional mismatch, enabling stable and beneficial reuse.
(a) Validation loss for $\sim\!450M$ parameter models
(b) Validation loss for $\sim\!1B$ parameter models
@@
Our primary contribution is ExoFormer, which resolves this tension by learning a set of dedicated exogenous anchor projections outside the sequential layer stack. The internal-anchor version, which we term NuResFormer (Normalized unified), serves as a comparative foundation. Decoupling of the two roles proves consistently beneficial: every ExoFormer variant outperforms its corresponding NuResFormer counterpart in perplexity while remaining competitive in downstream accuracy despite no increase in width or depth, with the dynamic variant achieving the best results. We explain this high performance via an Offloading Hypothesis: the external anchor assumes the role of preserving token identity, allowing the sequential layers to specialize almost exclusively in refinement.
@@ 2 Related Work
Zhou et al. (2025) introduced Value Residual Learning (ResFormer), which adds a residual connection from the first layer’s value projection ($V_{1}$) to the value projections of all subsequent layers (regulated by scalar coefficients $\lambda_{1}$ and $\lambda_{2}$). This simple yet effective method was shown to greatly improve model performance and data efficiency, highlighting the benefit of explicitly reusing early content representations. However, preliminary attempts to residualize queries and keys were found to be unstable (Zhou et al., 2025). Concurrently, gated attention mechanisms have been explored to introduce dynamic, input-dependent modulation to the attention output, improving expressiveness and training stability (Qiu et al., 2025). Our work generalizes residual learning to all attention pathways and stabilizes it via normalized mixing, bridging the gap between residual connections and gated attention.
@@
Recent work has sought to improve cross-layer information flow in Transformers beyond simple residual connections. Zhu et al. (2025) proposed Hyper-Connections that expand residual stream width. More recently, Xie et al. (2026) introduced Manifold-Constrained Hyper-Connections (mHC), restoring an identity-like signal-preservation property to the hyper-connected architecture.
@@
While techniques mentioned above and our work focus on enhancing residual pathways, they operate at different levels and can coexist. MUDDFormer primarily addresses the input to the attention projections. In contrast, our unified mixing framework focuses on mixing the attention projections themselves. Our approach can be seen as a form of residual blending after projection, whereas MUDDFormer enriches the input before projection.
@@ 3.1 Preliminaries and Notation
We consider three coefficient granularities for $\lambda$: Scalar (S) (a single coefficient shared across all channels), Headwise (H) (one coefficient per head, broadcast across $d_{k}$), and Elementwise (E) (one coefficient per channel).
@@ 3.2 Multi-Head Attention
$\begin{split}A_{n}&=\mathrm{softmax}\left(\frac{Q_{n}K_{n}^{\top}}{\sqrt{d_{k}}}\right)\in\mathbb{R}^{T\times T},\\ U_{n}&=A_{n}V_{n}\in\mathbb{R}^{T\times d_{\text{model}}},\end{split}$ (2)
@@
In NuResFormer/ExoFormer, the above stages use mixed tensors $\widehat{Q}_{n},\widehat{K}_{n},\widehat{V}_{n}$ from Eq. (10).
@@ 3.3 Gated Attention
In our attention block, we apply gating to the concatenated multi-head output (i.e., after Stage 2 and before Stage 3). Let $G_{n}\in\mathbb{R}^{T\times(hd_{k})}$ denote the gate logits, computed from $H_{n-1}$:
@@ 3.4 Unified Mixing Formulation
We now present a general framework for mixing. The core idea is to enrich each layer’s attention pathways with a set of persistent anchor projections $\{Q_{\text{anc}},K_{\text{anc}},V_{\text{anc}},G_{\text{anc}}\}$ that is reused, via learnable mixing, across all layers.
@@
The anchors are a fixed set of projections defined once for the entire model. In our work, we explore two instantiations:
@@
    $Q_{\text{anc}}$ $=H_{0}W^{Q}_{\text{anc}},$ $K_{\text{anc}}$ $=H_{0}W^{K}_{\text{anc}},$ (9)
    $V_{\text{anc}}$ $=H_{0}W^{V}_{\text{anc}},$ $G_{\text{anc}}$ $=H_{0}W^{G}_{\text{anc}},$
    where $W^{Q}_{\text{anc}},W^{K}_{\text{anc}},W^{V}_{\text{anc}},W^{G}_{\text{anc}}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are independent learnable weight matrices.
@@
$\widehat{S}_{n}$ $=\lambda^{S}_{n,1}\odot\mathrm{RMSNorm}(S_{\text{anc}})+\lambda^{S}_{n,2}\odot S_{n},$ (10)
$\qquad\forall S\in\{Q,K,V,G\},$
@@ 3.5 First Layer Tension and ExoFormer
While these roles appear misaligned—universal anchors favor invariance, while progressive computation necessitates change—they can theoretically coexist because it is an optional pathway modulated by learned mixing coefficients ($\lambda_{n,1},\lambda_{n,2}$). However, this forced coexistence places pressure the first layer to make compromises, inherently limiting its effectiveness in both roles. This tension is empirically supported by the analysis in Figure 7, showing that NuResFormer’s first layer adopts a permissive gating policy compared to standard baselines, indicating a compromise on selectivity in favor of serving as a stable anchor.
Our most successful approach, termed ExoFormer (where “Exo” signifies an external, outside-the-system anchor), instantiates the general framework of Section 3.4 with the dedicated exogenous projections of Eq. (9).
This decoupling proves beneficial: in our experiments, every ExoFormer variant outperformed its corresponding NuResFormer counterpart in perplexity, despite no change to depth or width.
@@ 3.6 Dynamic Mixing (DM) Module
Building upon the unified formulation, we take inspiration from MUDDFormer’s Depth-wise Aggregate (DA) module (Xiao et al., 2025) and introduce a dynamic variant where the learnable parameters are modulated by context-dependent scaling factors computed from the layer input $H_{n-1}$ using a small MLP. This allows the model to adapt its mixing strategy based on the specific context.
@@
For each component we compute using the dynamic scaling factors:
$\widehat{S}_{n}$ $=(\lambda^{S}_{n,1}\gamma^{S}_{n,1})\odot\text{RMSNorm}(S_{\text{anc}})+(\lambda^{S}_{n,2}\gamma^{S}_{n,2})\odot S_{n},$ (12)
$\qquad\forall S\in\{Q,K,V,G\},$
@@ 4.1 Experimental Setup
All models use a modern pre-normalized Transformer architecture with SwiGLU activations (Shazeer, 2020), QKNorm (Henry et al., 2020), and rotary position embeddings (Su et al., 2023). We follow prior work by initializing projection and classification layers to zero (Yang et al., 2022), removing bias terms (except for the dynamic mixing module) (Chowdhery et al., 2022), applying z-loss regularization (debrébisson2016zlossshiftscaleinvariant), and disabling dropout. Training is performed on the FineWeb-Edu dataset (Penedo et al., 2024).
Table 1: Performance comparison of model variants on 6 multiple-choice downstream tasks. The parameter counts are: 453M for models with gated attention, 454M for those without it, and 457M for models that incorporate an external anchor. Naïve Combination refers to the unmodified addition of Gated Attention and ResFormer. Dynamic ExoFormer achieves the highest overall accuracy and the lowest perplexity, demonstrating the advantages of decoupling the anchor from the first layer and using context-aware mixing. All models use full normalization on anchors unless noted otherwise. The notation V, G denotes ablations where only value and gate projections are mixed; queries and keys are computed solely from the current layer. All model labels refer specifically to Anchor Normalization (e.g., “Only Q/K Norm” means RMSNorm is applied only to anchor Q/K sources).
Model ARC-c ARC-e Hella. OBQA PIQA Wino. Avg. Acc PPL
Baselines
Base Transformer 30.97 63.38 42.02 33.60 67.30 51.54 48.14 14.79
Gated Attention 30.89 64.69 42.73 34.00 67.14 53.35 48.80 14.64
ResFormer (Value Residual) 33.62 64.86 43.49 34.40 68.88 52.64 49.65 14.32
Naïve Combination 32.94 64.52 43.47 32.60 68.34 52.64 49.09 14.25
Internal Anchor
E-NuResFormer (Only Q/K Norms) 31.74 64.27 43.83 33.00 68.23 52.41 48.91 14.21
H-NuResFormer (Only Q/K Norms) 34.73 64.06 44.45 34.20 67.57 53.04 49.68 14.21
S-NuResFormer (Only Q/K Norms) 32.85 66.08 43.87 34.20 68.28 53.67 49.83 14.22
E-NuResFormer (No Norm, {V,G}) 32.42 64.94 43.80 35.00 67.52 51.54 49.20 14.24
H-NuResFormer (No Norm, {V,G}) 31.23 64.18 43.50 34.00 67.68 54.06 49.11 14.22
S-NuResFormer (No Norm, {V,G}) 31.83 64.44 43.66 34.80 68.44 52.49 49.28 14.24
E-NuResFormer 33.70 65.07 44.23 33.60 68.34 53.12 49.68 14.15
H-NuResFormer 33.79 65.11 44.09 33.40 67.41 52.72 49.42 14.17
S-NuResFormer 32.42 64.02 43.93 33.20 67.90 53.20 49.11 14.17
External Anchor
E-ExoFormer (no norm) 32.08 63.76 43.49 34.40 68.77 55.64 49.69 14.30
Dynamic E-ExoFormer 33.36 65.87 44.54 34.40 68.28 55.17 50.27 14.09
E-ExoFormer 32.42 64.65 44.28 36.40 67.74 53.59 49.85 14.13
H-ExoFormer 32.17 65.87 44.00 32.40 68.23 52.72 49.23 14.14
S-ExoFormer 32.17 65.66 44.33 33.60 68.06 51.14 49.16 14.15
We optimize matrix parameters with Muon (specifically the Polar Express variant; Amsel et al. (2025)), applying a cautious weight decay of 0.1 (Chen et al., 2025), while 1D parameters are trained using AdamW without weight decay. Muon was utilized because it represents a state-of-the-art optimizer for large-scale LLM training (Liu et al., 2025). Gradient norms are clipped to 1.0, and all models follow the same optimization setup, to ensure fair comparison across architectures. Training uses a global batch size of 262,144 tokens and a sequence length of 2,048.
@@ 4.2 The Role of Anchor Normalization
RMSNorm serves two roles: first, acting as a scale-dependent rational function, it introduces a non-linear transformation into the otherwise linear residual pathway; and second, as a mild isotropization operator, RMSNorm projects representations onto the unit sphere while preserving directional information and removing scale differences. The unit vector retains the relative alignment between dimensions, which encodes semantic and syntactic features.
The necessity of applying RMSNorm to anchor sources is empirically supported by an analysis of the learned mixing coefficients in unnormalized models. For instance, in the ExoFormer variant without residual normalization, the proportion of near-zero coefficients (below 0.001) for $\lambda_{1}$ (the strength of anchor signal) is approximately triple that of the normalized variant. This suppression of the anchor pathway suggests the model is actively compensating for distributional mismatch.
We hypothesize that the naïve combination underperforms for the above reason. When applied to unbalanced mixtures, $\sigma(G)$ must compensate for distributional mismatch rather than execute content-aware filtering, resulting in instability during training (see Figure 2(a)) and worse downstream accuracy than ResFormer in isolation.
(a) Dynamic E-ExoFormer (full residual norms)
(b) E-ExoFormer (full residual norms)
Figure 3: Heatmaps showing the learned mixing coefficient ratio $\lambda_{n,1}/\lambda_{n,2}$ for each residualized component $\{Q,K,V,G\}$ across layers (y-axis) and channels/heads (x-axis) for model variants $\sim\!450M$. This ratio quantifies the model’s reliance on the anchor relative to the current layer’s projection; a higher value indicates stronger reuse of the early signal.
@@ 4.3 Extending Mixing to Q, K, and G Pathways
Consistent with prior observations (Zhou et al., 2025), we find a clear hierarchy of stability. Models attempting to use unnormalized Q/K residuals without QKNorm proved impossible to optimize, exhibiting divergent loss in preliminary runs. Critically, adding QKNorm alone stabilizes training and recovers baseline performance. We hypothesize this is because QKNorm compresses the scale of $Q$ and $K$, mitigating the distributional mismatch that arises when injecting early-layer routing signals into deeper layers. Furthermore, adding explicit RMSNorm to the Q/K anchor sources on top of QKNorm yields the best results, enabling positive reuse.